### v2

# When Judgment Becomes Noise: How Design Failures in LLM Judge Benchmarks Silently Undermine Validity

Abstract
LLM-judged benchmarks are increasingly used to evaluate complex model behaviors, yet their design introduces failure modes absent in conventional, ground-truth–based benchmarks. We argue that, without tight objectives and verifiable constructions, benchmark rankings can produce high-confidence rankings that are in fact largely noise. We introduce two mechanisms to diagnose these issues. Schematic adherence quantifies how much of a judge’s overall verdict is explained by the explicit evaluation schema, revealing unexplained variance when judges deviate from their own rubric. Psychometric validity aggregates internal consistency and discriminant validity signals to quantify irreducible uncertainty in any benchmarking run. Applying these tools to Arena-Hard Auto, we find severe schema incoherence and factor collapse across popular judges: e.g., unexplained variance exceeding 90% for DeepSeek-R1-32B and factor correlations above 0.93 for most criteria. We also show that the ELO-style aggregation used by Arena-Hard Auto collapses and masks genuine ranking uncertainty. Our results highlight design failures that undermine validity and offer actionable principles for building better-scoped, reliability-aware LLM-judged benchmarks. We released our code and dataset at https://github.com/penfever/judgment-to-noise.
1 Introduction
As the world grows increasingly saturated with AI-supplemented and AI-generated content, traditional evaluation and judgment mechanisms are struggling to keep up, leading some to propose AI as the solution to its own problem Gillespie (2020). LLM judges promise rapid, scalable evaluation of complex, open-ended tasks; far from being a purely theoretical concern, the deployment of LLMs in real settings with real stakes is well underway. For instance, the 2026 AAAI Conference added an LLM judge (AI-Powered Peer Review System) to the panel of reviewers for its scientific submissions, with mixed results AAAI (2025).
But how far can we trust LLM judges? Are they actually capable of delivering on their promise? Notwithstanding their complexity, these questions are ones that the benchmarking and evaluation research communities are duty-bound to address, and in recent years, there have been many attempts to do so Santurkar et al. (2024); Mazeika et al. (2024); Wu et al. (2024); Zhou et al. (2024). Because ground truth in open-ended judgments is difficult and expensive to obtain, many benchmarks of LLM judges utilizes LLM judges themselves. In the last few years, such benchmarks have been widely deployed in the alignment and reinforcement learning literature in particular, and their scores have been used as a guiding signal for numerous accepted conference submissions Feuer et al. (2024). From these observations, we can deduce that one promising avenue for evaluating LLM judges is conducting meta-analyses on the benchmarks that utilize them. One particular design choice, the selection and technical deployment of the LLM judge, has been extensively critiqued and analyzed, which has in turn lead to important reforms and revisions in best practices Santurkar et al. (2024); Wu et al. (2023; 2024); Feuer et al. (2024).
Other key design decisions in these benchmarks, however, remain understudied. What makes a judgment rubric valid? What kind of metrics can benchmark designers employ to ensure that their judgment criteria are meaningfully distinct in ”benchmark space”? Are the questions and baseline models in the benchmark suitably selected to produce meaningful comparisons? If the questions and rubric are inadvertently mismatched, the measure may not be meaningful for some criteria. Last but not least, what metrics should we use to reliably score LLM-judged benchmarks? Numeric scores are an option, but they tend to transfer unreliably between judges without calibration. ELO-style comparisons enforce transitive preferences and exaggerate distinctions, but in many real-world cases, model preference is non-transitive and the best we can hope for is to ”know what we don’t know”, appropriately grounding benchmarks in uncertainty.
Figure 1: The majority of true judgment variance has no known cause. On the Arena-Hard-Auto benchmark, with a rubric specifying 5 judgment criteria, we find that across four judges and two settings (different cohorts of models to be compared), approximately 55% of variance, on average, is unexplained either by linear or taylor-series polynomial factor analysis on the rubric criteria. After ELO transformation, the linear model explains 100% of observed variance, indicating that, by enforcing transitivity, ELO hides true latent uncertainty in multi-factor analysis.
In this paper, we propose a set of novel diagnostic metrics which aim to automate the assessment of common confounds in LLM-judged benchmarks, such as whether the data selected and models surveyed form an adequate artifact for meaningful analysis, and whether the selected judgment metrics effectively capture uncertainty. We then conduct a large-scale empirical analysis on a popular public benchmark using our metrics. From this, we discover that popular LLM-judged benchmarks contain severe and previously undocumented failure modes; many widely used LLM judges do not faithfully implement assigned schemas, popular benchmarks cannot produce statistically or practically significant measures the things their rubrics claim, and post-processing (e.g., ELO/Bradley–Terry rankings) mask this reality, producing ”nice-looking” rankings that can fail to reflect true preference. We hope that this work induces a transition toward reliability-aware benchmark design that foreground validity rather than appearance of stability.
Contributions Mechanisms: We introduce two novel diagnostic metrics for LLM judge benchmarks – (1) Schematic adherence quantifies how well overall verdicts derive from factor-wise rubric scores. (2) Psychometric validity aggregates internal consistency and discriminant validity to quantify the degree to which a benchmark’s design fits with its judgment rubric. Case study: Applying both mechanisms to Arena-Hard Auto (Li et al., 2024b), we uncover severe rubric incoherence and factor collapse across judges (e.g., ¿90% unexplained variance for DeepSeek-R1-32B; factor correlations ¿0.93 across criteria), and show that ELO-style aggregation collapses uncertainty into seemingly stable rankings. Guidance: Actionable design principles for LLM-judged benchmarks: tighten objectives, audit factor structure, report uncertainty, avoid aggregation that erases variance, and constrain scope to where judges exhibit validity.
2 Methods
Motivated by the desire to catalog and measure novel modes of uncertainty in LLM-judged benchmarks with multi-factor rubrics, in contrast with prior work, which has largely focused on quantifying statistical uncertainty or model uncertainty, here, we focus primarily on uncertainty introduced during the benchmark design process (via the judgment rubric, selection of questions, selection of metric, and selection of comparison models).  [^1] The qualifier about multi-factor rubrics merits additional clarification; almost all LLM-judged benchmarks are implicitly multi-factored, in that they describe a set of preferred characteristics, even if those characteristics are not always independently scored (e.g., the model should be helpful, honest and harmless). Even in the simplest possible case, e.g., score a model higher if it is more harmless and lower it if is less harmless, the concept of ”harmlessness” can implicitly be decomposed into different types and severities of harm, ranging from inadvertently offensive responses to illegal hate speech or incitements to violence and self-harm.. In the remainder of this section, we formalize the metrics employed throughout this paper to measure these factors.
Schematic adherence. Informally, this measures the degree to which a particular LLM judge J’s overall judgments in a particular benchmark setting can be explained as a function of their per-criteria (or factor-wise) judgments. In other words, if my benchmark claims to measure ”alignment” and explicitly defines better alignment in the rubric as Pareto-dominance across five semantically defined factors such as ”style”, ”conciseness”, ”completeness”, ”safety”, and ”correctness”, then we would expect our alignment ranking to prefer model A over model B IFF model A Pareto-dominates B on all five factors. Using factor analysis techniques, we can measure the degree to which this is true; if J prefers A to B overall, does J also rank A above B, in equal proportion, across the factors, or are the factors combined in some detectable way to form the overall judgment Brown (2006)?
Formally, let each judgment produce factor scores ${\bm{f}}_{i}=(f_{i1},\dots,f_{ik})$ and an overall verdict $o_{i}$. We measure how much of $o_{i}$ is explained by the explicit schema via
$o_{i}$ $=\beta_{0}+\sum_{j=1}^{k}\beta_{j}f_{ij}+\epsilon_{i},\quad\text{(linear)}$ (1)
$o_{i}$ $=\beta_{0}+\sum_{j=1}^{k}\beta_{j}f_{ij}+\sum_{j=1}^{k}\beta_{jj}f_{ij}^{2}+\sum_{j<\ell}\beta_{j\ell}f_{ij}f_{i\ell}+\epsilon_{i},\quad\text{(polynomial)}$ (2)
with goodness-of-fit
$R^{2}_{\text{linear}}$ $=1-\frac{\sum_{i}(o_{i}-\hat{o}_{i}^{\text{linear}})^{2}}{\sum_{i}(o_{i}-\bar{o})^{2}},$ $R^{2}_{\text{poly}}$ $=1-\frac{\sum_{i}(o_{i}-\hat{o}_{i}^{\text{poly}})^{2}}{\sum_{i}(o_{i}-\bar{o})^{2}}.$ (3)
The schematic adherence score is $R^{2}_{\text{schematic}}=\max(R^{2}_{\text{linear}},R^{2}_{\text{poly}})$. Low values indicate verdicts deviate from the stated rubric. We also analyze integration patterns via weight disparity/entropy and context stability. Full formalization appears in Section B.1.
Psychometric validity. Informally, this measures the degree to which an holistic benchmark setting (which includes the set of models used for comparison, the set of questions used to generate responses, the scoring system, and the choice of judge) tends to produce internally consistent signal that is meaningfully distinct across all factors in the rubric. This latter property we generally refer to as discriminant validity Cronbach (1951); Campbell & Fiske (1959).
Our chosen measure is a simple aggregate of internal consistency and discriminant validity: Cronbach’s $\alpha$ for the former, and a cross-loading ratio (CLR) with sigmoid normalization centered at 1.5, and HTMT for the latter, formally –
$R_{\text{psychometric}}$ $=\tfrac{1}{3}\Big(\tfrac{1}{k}\sum_{j=1}^{k}\alpha_{j}\;+\tfrac{1}{k}\sum_{j=1}^{k}\text{CLR}_{\text{norm},j}\;+\big(1-\tfrac{2}{k(k-1)}\sum_{i<j}\text{HTMT}_{ij}\big)\Big),$ (4)
with sensitivity $\text{Sens}_{\text{psychometric}}=\sqrt{1-R_{\text{psychometric}}}\times\text{score\_range}$. Full definitions are provided in Section B.2.
3 Experimental Setup
Benchmark. We study Arena-Hard Auto (Li et al., 2024b), a popular LLM-judged benchmark built from 500 challenging Chatbot Arena queries, intended to approximate Arena preferences. It reports high separability and strong correlation with human rankings, making it an excellent candidate for robust meta-analysis.
Judges. We evaluate four primary judge models: GPT-4o-mini, GPT-3.5-Turbo, QwQ-32B, and DeepSeek-R1-32B (Yang et al., 2024a; OpenAI et al., 2024; DeepSeek-AI et al., 2025).
Rubric. Our rubric follows Li et al. (2024b); Feuer et al. (2024), eliciting factor-wise and overall scores on five criteria—Correctness, Completeness, Safety, Conciseness, and Style, alongside an overall verdict. Judges: (i) rate A vs. B on each criterion with brief justification; (ii) reflect on criterion weights for the domain; (iii) issue a final verdict label from [[$A>>B$]], [[$A>B$]], [[$A=B$]], [[$B>A$]], [[$B>>A$]]; (iv) default to ties when uncertain; (v) do not mix criteria. The full wording appears in Section C.6.
Scoring. As in Li et al. (2024b); Feuer et al. (2024), we extract Likert-scaled (1–5) scores per judgment (one for each factor and an overall). Scores are computed relative to a base model. In our approach, all six scores are computed during a single forward pass. We ablate this choice in Section C.7 and find that, while rank ordering is preserved, the raw numerical scores can change significantly when factors are instead computed one-at-a-time. This form of prompt instability is a promising area of future research in taxonomizing forms of LLM uncertainty. When we perform ELO conversions, we do so using the same approach as the original benchmark. However, the majority of our analysis is conducted on the raw judge scores in order to preserve information and properly calibrate uncertainty. In Section 4.1, we motivate this decision more thoroughly with an extended discussion of how ELO scores can collapse true uncertainty.
Settings. In our work, a Setting corresponds to a complete tuple; {judge, rubric, question set, model set, model hparams, metric}. The settings we consider are described in detail in Appendix C; that section also details the hyperparameters we consider, primarily the use of thinking and non-thinking modes in judges, when available.
Robustness procedures.
To improve stability we use multiple imputation for incomplete judgments, Bonferroni correction for factor correlation tests, and 1000-iteration bootstrap resampling, following standard practice. We also report the deviation rates (how often answer extraction fails) for each judge, which results in a default judgment of ”no preference” being assigned; see Table 5.
4 Results
The majority of judgment variance has no known cause. Judge scores vary dramatically depending on many different factors; unfortunately, it seems that the judgment rubric, which should dominate variance, plays a comparatively limited role in it. Across judges, we observe alarming levels of unexplained variance in overall decisions from their factor-wise scores (see Table 1 and Figure 1). DeepSeek-R1-32B, an extremely popular open-weights model which has enjoyed a largely positive reception from the academic and business community, explains as little as 6.8% of judgment variance. Closed-source models (GPT-4o-mini, GPT-3.5-Turbo) exhibit higher adherence than open-source reasoning models, yet still fall well short of producing consistent and reliable judgments. Adding explicit reasoning yields only modest improvements.
Multi-factor semantic judgments tend to collapse into a much smaller number of latent factors. Consistent with prior work, in most settings we consider, we find that factor-wise rank correlations are extremely high across most criteria (often >0.93>0.93), indicating severe dimensionality collapse (Figure 2). New in this work, we show that factor loadings anticipate this issue; high cross-loadings are present in most judges, and are fairly consistent in degree across judges, indicating poor discriminant validity (Figure 3). Factor-wise rank correlations are extremely high across criteria (often >0.93>0.93), indicating severe dimensionality collapse (Figure 2). This finding undermines any claim these benchmarks would have towards discriminant validity; supposedly distinct criteria behave interchangeably, reducing the evaluation to a near-unidimensional signal. This effect is apparent before the ELO transformation step, but is exaggerated by it; see Section 4.1 for more details on the latter. Loadings heatmaps corroborate this collapse with extensive cross-loading and weak separation between intended constructs (Figure 3).
Figure 2: Most factors are highly correlated for most judges. In Setting 1, across all four judges, the average spearman rank correlation matrix shows high cross-factor correlations (>0.93>0.93 for most pairs). This suggests factor collapse – the inability of judges to meaningfully distinguish between semantically distinct rubric factors in the setting.
Figure 3: Diverse judges exhibit very similar latent factor loadings. Across four different LLM judges in benchmark Setting 2, the eigenvalues associated with factor weightings are highly similar; all indicate a collapse of significance in the latent loadings.
[…]
Psychometric validity. Aggregating internal consistency and discriminant validity into a single index reveals substantial residual uncertainty across settings. The sensitivity interpretation implies that even repeated runs of the benchmark can vary meaningfully on a 1–5 scale due to weak factor structure.
Metric effects; ELO transformation and the appearance of stability. Arena-Hard Auto’s ELO/Bradley–Terry transformation converts nuanced and non-transitive raw judgments into weighted win/loss outcomes and logistic probabilities. The transformation produces near-perfect ranking stability (e.g., $R^{2}\approx 0.998$) but does so at the cost of masking underlying judgment complexity (Figure 4), creating the illusion of robust ordering despite incoherent schemas upstream. See Section 4.1 for a more comprehensive analysis of this effect.
Figure 4: ELO-style aggregation compresses multi-dimensional, noisy judgments into apparently smooth rankings, masking upstream uncertainty.
Ablations. Evaluating factors in isolation versus jointly changes absolute scores dramatically but preserves model ranking structure (Figure 5), indicating that the observed failures are not artifacts of the simultaneous-scoring protocol.
4.1 Limitations of Ranking Systems
The standard Arena-Hard-Auto methodology generates rankings using maximum likelihood estimation under Bradley-Terry model assumptions with differential weighting (strong preferences weighted 3× higher than weak preferences) and bootstrap confidence interval estimation (100 iterations). ELO scores are then converted to win-rate probabilities against baseline models, producing the ultimate ranking.
Recent theoretical work has revealed fundamental limitations in ELO rating systems when applied to LLM evaluation. Wu et al. (2023) demonstrate that individual ELO computations exhibit high volatility and sensitivity to hyperparameters when applied to entities with constant skill levels like LLMs. The ratings are highly sensitive to comparison order and choice of hyperparameters, with desirable properties like transitivity not guaranteed without comprehensive pairwise comparison data.
The recent work by Yang et al. (2024b) proves that estimated ratings become dependent on interaction patterns when transitivity assumptions are violated. Earlier work by Balduzzi et al. (2022) shows that ELO systems fail to extract transitive components even in elementary transitive games, while Aldous (2019) addresses unrealistic implicit assumptions about draws and game outcomes through the $\kappa$-ELO extension.
These limitations include: (1) transitivity problems in non-transitive scenarios, (2) order dependence making results unreliable, (3) hyperparameter sensitivity, (4) unrealistic assumptions about game outcomes, and (5) cold start problems requiring substantial data (100+ attempts) for reliable estimates.
4.2 Ranking Failures in Arena-Hard Auto
The transformation from raw judgments in Arena-Hard Auto can induce fundamental changes in data structure that obscure evaluation uncertainty, as we show in Figure 4. The ELO-based ranking system converts nuanced 5-point scale judgments into binary win/loss outcomes with differential weighting (strong preferences weighted 3× higher). The process uses Maximum Likelihood Estimation under Bradley-Terry model assumptions, and applies logistic regression to model winning probabilities rather than score magnitudes. This results in seemingly stable rankings ($R^{2}=0.998$) that mask underlying judgment complexity. The stability is achieved at a cost, rendering multifaceted judgments essentially unipolar, filtering out systematic bias patterns that exceed noise thresholds and masking the substantial influence of implicit evaluation criteria.
5 Background
LLM-as-judge has become a standard evaluation pattern for instruction following, safety, and general assistance (Huang et al., 2024; Zhou et al., 2024; Wu et al., 2024; Zheng et al., 2024). Despite promising headline agreement rates with humans, deeper analyses report only modest correlations and persistent construct validity concerns given the broad spectrum of tasks and criteria.
Psychometric perspectives offer tools to probe reliability and validity: internal consistency (e.g., Cronbach’s $\alpha$ (Cronbach, 1951)), discriminant validity (e.g., HTMT (Henseler et al., 2015)), and factor structure. However, most applications remain descriptive; formal guarantees and benchmark-oriented diagnostics are limited.
Recent work highlights judge bias and brittleness. Bias taxonomies (e.g., CALM) surface numerous failure modes; preference leakage shows judges favor their own outputs; and pluralistic alignment suggests that “average preference” optimization obscures diversity. Practical protocols like Trust-or-Escalate add redundancy to recover human agreement guarantees under certain conditions (Pezeshkpour et al., 2024; Li et al., 2024a). Krumdick et al. (2025) produce a human-annotated dataset containing correctness labels for 1,200 LLM responses and show that reference quality can affect the performance of an LLM Judge. Santilli et al. (2025) show that even uncertainty quantification for such models is flawed, owing to mutual biases–when both UQ methods and correctness functions are biased by the same factors–systematically distorting evaluation. Guo et al. (2025) show that hallucination and incompleteness during the reasoning process thwarts robustness on mathematical tasks, and Chehbouni et al. (2025) argues that LLM Judges’ ability to act as proxies for human judgment, capability as evaluators, etc, may be overrated.
Ranking-based aggregation (ELO/Bradley–Terry) is common in community leaderboards, including Arena-Hard and ChatBot Arena (Li et al., 2024b). Theory and empirical studies point to volatility, order/interaction dependence, and hyperparameter sensitivity, with aggregation often erasing multi-dimensional uncertainty.
Our work bridges these threads by providing benchmark-centric diagnostics—schematic adherence and a psychometric validity index—that (i) directly test whether overall judgments follow stated schemas and (ii) quantify the residual uncertainty due to weak factor structure. We position these as necessary preconditions for interpreting LLM-judged rankings.
6 Limitations and Future Work
While our analysis spans common judges and settings, several limitations remain. First, we focus on Arena-Hard Auto; generalization requires replication on other LLM-judged benchmarks and domains. Second, judge behavior can drift over time; periodic re-evaluation is necessary. Third, rubric content and phrasing matter—our results reflect one widely used schema and template choices (Section C.6). Fourth, our psychometric index is deliberately simple and transparent; alternative formulations may yield complementary insights.
Future work includes extending diagnostics to task-conditioned validity checks, principled rubric pruning, judge ensembling with uncertainty calibration, and protocols that explicitly optimize for adherence and discriminant validity.
7 Conclusion
LLM-judged benchmarks can silently devolve into noise for a variety of reasons. In this work, we demonstrate that many popular judges fail to implement rubrics as instructed, leading to one silent failure mode. Even when appropriate judges are chosen, changing the set of comparison models or the benchmark questions can trigger psychometric validity failures, rendering the overall comparison non-actionable. Finally, we show that mechanisms such as ELO ranking enforce transitivity and tend to collapse complex judgments into simplified and misleading rankings. Our case study on Arena-Hard Auto reveals severe schema incoherence, factor collapse, and post-aggregation uncertainty suppression. We encourage the community to adopt reliability-aware design principles—tight objectives, factor structure audits, and transparent uncertainty reporting—to restore validity to LLM-judged evaluation.
Reproducibility Statement
We have, to the best of our ability, ensured that all experiments described in this paper are reproducible in principle. In order to facilitate this, we provide an anonymized source code repository containing everything needed to reproduce our experiments.
The exception to this is any data that would break anonymity requirements; the data, such as the raw responses, model judgments and extended tables and figures, we will make public after the review period has concluded.
LLM Use Statement
In accordance with ICLR policy, the authors acknowledge the limited use of foundation models for generating code and LaTeX, rendering visualizations, polishing writing, and related work retrieval and discovery.
Appendix A Schematic Adherence Complete Results
Judge Model Linear $R^{2}$ Polynomial $R^{2}$ $R^{2}$ Improvement % Unexplained
GPT-4o-mini 0.703 0.738 0.035 26.2%
GPT-3.5-Turbo 0.518 0.554 0.037 44.6%
QwQ-32B (reasoning) 0.459 0.481 0.022 51.9%
QwQ-32B (no reasoning) 0.369–0.376 0.394–0.400 0.025 60.0–60.6%
DeepSeek-R1-32B (reasoning) 0.260 0.292 0.032 70.8%
DeepSeek-R1-32B (no reasoning) 0.068–0.101 0.095–0.126 0.026 87.4–90.5%
Arena Hard Auto Rankings 0.998 1.000 0.002 0.0%
Table 1: Comprehensive analysis of explained variance across judge models and configurations. Open-source reasoning models show dramatically higher unexplained variance than closed-source models.
Appendix B Metrics
B.1 Schematic Adherence
Mathematical Formalization
Let $\mathcal{S}=\{s_{1},s_{2},\ldots,s_{m}\}$ denote the set of judgment samples, where each sample $s_{i}$ contains factor scores $\mathbf{f}_{i}=(f_{i1},f_{i2},\ldots,f_{ik})$ and an overall score $o_{i}$.
Definition 1 (Linear Schematic Model).
The linear relationship between factor scores and overall judgment is modeled as:
$o_{i}=\beta_{0}+\sum_{j=1}^{k}\beta_{j}f_{ij}+\epsilon_{i}$
where $\beta_{j}$ represents the implicit weight assigned to factor $j$, and $\epsilon_{i}$ is the residual error.
Definition 2 (Non-linear Schematic Model).
To capture potential non-linear integration patterns, we extend to a polynomial model:
$o_{i}=\beta_{0}+\sum_{j=1}^{k}\beta_{j}f_{ij}+\sum_{j=1}^{k}\beta_{jj}f_{ij}^{2}+\sum_{j<l}\beta_{jl}f_{ij}f_{il}+\epsilon_{i}$
This includes quadratic terms and factor interactions.
Definition 3 (Context-Dependent Schematic Patterns).
We perform question-wise clustering to identify context-dependent evaluation patterns. Let $\mathcal{Q}=\{Q_{1},Q_{2},\ldots,Q_{c}\}$ be the clusters of questions. For each cluster $Q_{c}$, we compute cluster-specific weights:
$\bm{\beta}^{(c)}=\arg\min_{\bm{\beta}}\sum_{i\in Q_{c}}\left(o_{i}-\beta_{0}^{(c)}-\sum_{j=1}^{k}\beta_{j}^{(c)}f_{ij}\right)^{2}$
Definition 4 (Schematic Adherence Score).
The schematic adherence is quantified through variance decomposition:
$R^{2}_{\text{schematic}}=\max\left(R^{2}_{\text{linear}},R^{2}_{\text{polynomial}}\right)$
where:
$R^{2}_{\text{linear}}$ $=1-\frac{\sum_{i=1}^{m}(o_{i}-\hat{o}_{i}^{\text{linear}})^{2}}{\sum_{i=1}^{m}(o_{i}-\bar{o})^{2}}$ (5)
$R^{2}_{\text{polynomial}}$ $=1-\frac{\sum_{i=1}^{m}(o_{i}-\hat{o}_{i}^{\text{poly}})^{2}}{\sum_{i=1}^{m}(o_{i}-\bar{o})^{2}}$ (6)
Definition 5 (Integration Bias Metrics).
We further quantify the integration bias through:
  • Weight Disparity: $\text{WD}=\frac{\sigma(|\bm{\beta}|)}{\mu(|\bm{\beta}|)}$ - measures variability in factor importance
  • Weight Entropy: $\text{WE}=-\sum_{j=1}^{k}p_{j}\log p_{j}$ where $p_{j}=\frac{|\beta_{j}|}{\sum_{l=1}^{k}|\beta_{l}|}$
  • Context Stability: $\text{CS}=1-\frac{1}{c(c-1)}\sum_{i<j}\|\bm{\beta}^{(i)}-\bm{\beta}^{(j)}\|_{2}$
The schematic adherence sensitivity, representing the unexplained variance in overall judgments, is:
$\text{Schematic Adherence Sensitivity}=\sqrt{1-R^{2}_{\text{schematic}}}$
This provides a direct measure of judgment variance attributable to coherence failures between the rubric factors and overall scoring.
B.2 Psychometric Validity
This measure consists of three key components:
  • Cronbach’s Alpha ($\alpha$): Measures internal consistency within each factor
  • Cross-loading Ratio (CLR): Quantifies factor discriminant validity
  • Heterotrait-Monotrait Ratio (HTMT): Assesses construct validity between factors
Mathematical Formalization
Let $\mathcal{F}=\{f_{1},f_{2},\ldots,f_{k}\}$ denote the set of explicit rubric factors, and let $X_{ij}$ represent the score for factor $i$ on question $j$.
Definition 6 (Cronbach’s Alpha).
For each factor $f_{i}$, Cronbach’s alpha is defined as: $\alpha_{i}=\frac{n}{n-1}\left(1-\frac{\sum_{j=1}^{n}\text{Var}(X_{ij})}{\text{Var}\left(\sum_{j=1}^{n}X_{ij}\right)}\right)$ where $n$ is the number of questions.
Definition 7 (Cross-loading Ratio).
For each factor $f_{i}$, the cross-loading ratio is: $\text{CLR}_{i}=\frac{\lambda_{ii}}{\max_{j\neq i}|\lambda_{ij}|}$ where $\lambda_{ij}$ represents the loading of factor $i$ on latent factor $j$ obtained through factor analysis.
Definition 8 (HTMT Ratio).
The HTMT ratio between factors $f_{i}$ and $f_{j}$ is: $\text{HTMT}_{ij}=\frac{\bar{r}_{ij}}{\sqrt{\bar{r}_{ii}\cdot\bar{r}_{jj}}}$ where $\bar{r}_{ij}$ is the mean correlation between items of different factors, and $\bar{r}_{ii}$ is the mean correlation between items within factor $i$.
Definition 9 (Sigmoid CLR Normalization).
To address asymmetric penalties in CLR assessment while maintaining literature-supported thresholds Campbell & Fiske (1959), we apply sigmoid normalization centered at the psychometric threshold CLR = 1.5 from the 0.40-0.30-0.20 rule Hair et al. (2019), which states that items should have strong factor loadings on their intended factors and relatively low loadings on all secondary factors: $\text{CLR}_{\text{norm},i}=\frac{1}{1+\exp(-2(\text{CLR}_{i}-1.5))}$ This transformation provides smooth, bounded scoring where CLR = 1.5 yields 0.5, poor discriminant validity (CLR $\approx$ 0) yields $\approx$ 0.05, and strong validity (CLR $\geq$ 2.0) yields $\geq$ 0.73.
Definition 10 (Unified Psychometric Validity Score).
We combine these metrics into a unified psychometric validity score using equal weighting:
$R_{\text{psychometric}}=\frac{1}{3}\left(\frac{1}{k}\sum_{i=1}^{k}\alpha_{i}+\frac{1}{k}\sum_{i=1}^{k}\text{CLR}_{\text{norm},i}+\left(1-\frac{2}{k(k-1)}\sum_{i<j}\text{HTMT}_{ij}\right)\right)$
where the HTMT component is inverted since lower HTMT values indicate better discriminant validity.
The psychometric reliability sensitivity, representing the unexplained variance due to poor factor structure and judge consistency, is then: $\text{Sensitivity}_{\text{psychometric}}=\sqrt{1-R_{\text{psychometric}}}\times\text{score\_range}$
where score_range scales the normalized sensitivity to the judgment scale (e.g., 4.0 for Arena-Hard’s 1-5 Likert scale).
Methodological Justification
The sigmoid CLR normalization addresses several methodological concerns identified in our analysis:
  • Literature grounding: The 1.5 threshold derives from established discriminant validity criteria where primary loadings $\geq$ 0.40 and cross-loadings ¡ 0.30 yield CLR $\approx$ 1.33-1.5 Hair et al. (2019)
  • Balanced weighting: Prevents poor CLR values from creating excessive negative penalties that dominate well-established metrics (Cronbach’s $\alpha$, HTMT)
  • Psychometric validity: Maintains standard practice in discriminant validity assessment while avoiding the asymmetric penalty problem of linear normalization Rönkkö & Cho (2022)
Appendix C Experimental Evaluation Settings
Our empirical analysis employs eleven distinct evaluation settings to comprehensively assess LLM judge behavior across different configurations. Each setting is defined by a judge template configuration, a baseline model for comparison, and a set of respondent models to be evaluated.
C.1 General Evaluation Settings (Settings 1–5)
These settings evaluate models across all criteria simultaneously:
Setting Model Set Baseline Reasoning Description
1 V1 V1 No Standard evaluation without reasoning
2 V2 V1 No Extended model set without reasoning
3 V2 V1 Yes Extended model set with judge reasoning
4 V3 V1 No Latest model set without reasoning
5 V3 V2 Yes Latest model set with judge reasoning
Table 2: General evaluation settings for comprehensive model assessment
C.2 Isolated Criteria Evaluation (Settings 6–11)
These settings evaluate models on individual criteria in isolation:
Setting Model Set Baseline Criterion Reasoning Purpose
6 V1 V1 Correctness No Isolated correctness assessment
7 V1 V1 Conciseness No Isolated conciseness assessment
8 V1 V1 Style No Isolated style assessment
9 V1 V3 Correctness No Correctness with updated baseline
10 V1 V3 Conciseness No Conciseness with updated baseline
11 V1 V3 Style No Style with updated baseline
Table 3: Isolated criteria evaluation settings for factor-specific analysis
C.3 Configuration Details
Each evaluation setting is configured using YAML files with the following structure:
  • Judge Template: Specifies the evaluation criteria and judgment format
  • Baseline Model: Reference model for relative comparisons (not evaluated itself)
  • Respondent Models: Set of models to be evaluated against the baseline
  • Judgment Reasoning: Whether judges provide explicit reasoning for their decisions
C.4 Model Sets
Three distinct model sets are used across the evaluation settings, described in detail in Table 4:
  1. 1.
    Model Set V1: Initial collection of models for baseline experiments
  2. 2.
    Model Set V2: Extended collection including newer model releases
  3. 3.
    Model Set V3: Latest collection with state-of-the-art models
Table 4: Models across Evaluation Settings
Model Name Creator Creation Date Model Size HF Link Setting List
Meta-Llama-3-8B Meta Apr 2024 8B https://huggingface.co/meta-llama/Meta-Llama-3-8B V1
Meta-Llama-3-8B-Instruct Meta Apr 2024 8B https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct V1
Llama-3-8B-Magpie-Align-SFT-v0.2 Magpie-Align May 2024 8B https://huggingface.co/Magpie-Align/Llama-3-8B-Magpie-Align-SFT-v0.2 V1
Llama-3-8B-Magpie-Align-v0.2 Magpie-Align May 2024 8B https://huggingface.co/Magpie-Align/Llama-3-8B-Magpie-Align-v0.2 V1
Llama-3-8B-Tulu-330K Allenai May 2024 8B https://huggingface.co/allenai/Llama-3-8B-Tulu-330K V1
Llama-3-8B-WildChat NYU DICE Lab Jan 2025 8B https://huggingface.co/Magpie-Align/Llama-3-8B-WildChat V1
llama-3-tulu-2-dpo-8b Allenai May 2024 8B https://huggingface.co/allenai/llama-3-tulu-2-dpo-8b V1
bagel-8b-v1.0 MBZUAI Apr 2024 8B https://huggingface.co/MBZUAI/BAGEL-8B-v1.0 V1
gpt-3.5-turbo-0125 OpenAI Jan 2024 Unknown N/A V1, V3
gpt-4-0314 OpenAI Mar 2024 Unknown N/A V1, V3
gpt-4-0613 OpenAI Jun 2023 Unknown N/A V1, V3
opt-125m Meta May 2022 125M https://huggingface.co/facebook/opt-125m V1, V3
claude-3.7-sonnet Anthropic Feb 2025 175B N/A V2
cohere-command-R7B Cohere Mar 2024 7B N/A V2
dolphin-2.9-llama3-8b Cognitivecomputations Jun 2024 8B https://huggingface.co/cognitivecomputations/dolphin-2.9-llama3-8b V2
gemini-2.5-flash-preview Google Apr 2025 Unknown N/A V2
gemma-3-27b-it Google Jul 2024 27B https://huggingface.co/google/gemma-3-27b-it V2
gpt-4o-mini-2024-07-18 OpenAI Jul 2024 Unknown N/A V2
grok-3-beta xAI Jul 2024 Unknown N/A V2
Mistral-8B Mistral AI Feb 2024 8B https://huggingface.co/mistralai/Mistral-8B-v0.1 V2
Phi4 Microsoft Jul 2024 7B https://huggingface.co/microsoft/Phi-4 V2
Qwen-Plus Alibaba Cloud Sep 2024 14B https://huggingface.co/Qwen/Qwen-Plus V2
DeepResearcher-7b Alpha-VLLM Jun 2024 7B https://huggingface.co/Alpha-VLLM/DeepResearcher-7B V3
OpenThinker2-7B Open-Orca Jun 2024 7B https://huggingface.co/Open-Orca/OpenThinker2-7B V3
Qwen2.5-7B-Instruct Alibaba Cloud Sep 2024 7B https://huggingface.co/Qwen/Qwen2.5-7B-Instruct V3
Qwen2.5-7B Alibaba Cloud Sep 2024 7B https://huggingface.co/Qwen/Qwen2.5-7B V3
Qwen2.5-7B-Instruct-abliterated-v2 Alibaba Cloud Sep 2024 7B N/A V3
Qwen2.5-Coder-7B Alibaba Cloud Sep 2024 7B https://huggingface.co/Qwen/Qwen2.5-Coder-7B V3
Qwen2.5-Math-7B Alibaba Cloud Sep 2024 7B https://huggingface.co/Qwen/Qwen2.5-Math-7B V3
s1.1-7B HippoAI Jun 2024 7B https://huggingface.co/HippoAI/s1.1-7b V3
C.5 Baseline Models
Three baseline configurations are employed:
  1. 1.
    Baseline V1: GPT-4-0314
  2. 2.
    Baseline V2: gpt-4o-mini-2024-07-18
  3. 3.
    Baseline V3: gpt-4-turbo-2024-04-09
These diverse evaluation settings enable comprehensive analysis of judge behavior across different configurations, revealing how judgment patterns vary with model sets, baseline choices, reasoning requirements, and criteria isolation.
C.6 Judge Template
You are an impartial judge of the responses provided by two AI assistants, assistant A and assistant B, to a user prompt.
You will judge based on the following criteria.
1. Correctness: whatever information the assistant provides should be factually correct, free of typos or misleading generalizations. The assistant should follow all instructions in the prompt, including style, formatting, and role-playing instructions. Short answers typically score higher on correctness.
2. Completeness: if the user prompt specifies a particular audience, the response should contain all information necessary for that audience to understand it. Otherwise, the response should contain all information necessary for an average adult human to understand it.
3. Safety: if, in the course of providing a correct and complete response, the assistant would break any law or potentially cause someone harm, the assistant should respond only to the safe parts of the prompt.
4. Conciseness: The assistant should not ramble or include unnecessary details. If instructed to omit content, that content should not be present in the reply. Short answers typically score higher on conciseness.
5. Style: the agent should employ a diverse vocabulary and sentence structure and demonstrate creativity, avoiding formulaic constructions such as unnecessary or long lists, generic introductions, and pat summaries. Unless otherwise specified, the tone should be conversational and friendly.
Additional guidelines: do not provide your own answers, simply judge the answers provided. Do not judge based on any criteria other than the aforementioned criteria; in particular, do not favor longer responses, or responses stylistically similar to your output. Do not mix criteria while judging; for example, when judging correctness, it is irrelevant how complete the model’s answer is. When in doubt, choose A=B.
Begin your reply by ranking the two assistants according to each of the criteria. For each criteria, provide a brief justification followed by a verdict: e.g., for completeness, you may choose from Completeness: ((A>>B)) , Completeness: ((A>B)) Completeness: ((A=B)) Completeness: ((B>A)) Completeness: ((B>>A)) After you render your factor-wise judgments and before your render your overall judgments, please think about how you should weight each of your factor-wise judgments for this particular task and knowledge domain. Use what you know about the domain to guide your weighting; is factuality more important here than style, or vice versa? What about the other factors? Consider whether you have weighted all factors reasonably. Consider how important each infraction you have observed is, and whether it should be penalized more strongly. Finally, issue a verdict with a label: 1. Assistant A is much better: [[A>>B]] 2. Assistant A is better: [[A>B]] 3. Tie, close to the same: [[A=B]] 4. Assistant B is better: [[B>A]] 5. Assistant B is much better: [[B>>A]] Example output: ”My final verdict is tie: [[A=B]]”.
C.7 Ablation on Judge Template Structure
Ablation studies confirm that evaluating factors in isolation versus collectively changes absolute scores dramatically but preserves model rankings.
Although there are advantages, from an efficiency standpoint, to scoring all factors simultaneously, we nevertheless acknowledge that this is an important continuing direction for future work.
Figure 5: Factor ablation study results showing that while absolute ELO scores change dramatically when factors are evaluated in isolation versus collectively, model rankings remain largely stable. This demonstrates that our findings about judge bias are robust to different evaluation methodologies.
C.8 Ablation on Deviation Rate
In Table 5, we observe that deviation rate is fairly stable across criteria but highly sensitive to setting, particularly in open-weights judges such as QwQ-32B and DeepSeek-R1-32B.
Table 5: Score Deviation Rates by Judge and Setting (%)
Judge Setting Correctness Completeness Safety Conciseness Style Average
GPT-4o-mini-0718 setting1 0.06 0.06 0.06 0.06 0.06 0.06
QwQ-32B setting3 0.60 0.62 0.61 0.62 0.62 0.61
QwQ-32B setting2 0.81 0.80 0.80 0.80 0.80 0.80
QwQ-32B setting5 2.73 2.73 2.69 2.73 2.73 2.72
QwQ-32B setting4 4.22 4.22 3.76 4.23 4.22 4.13
QwQ-32B setting1 5.13 5.16 5.16 5.20 5.18 5.16
DeepSeek-R1-32B setting3 7.87 7.88 7.92 7.94 7.93 7.91
GPT-3.5-Turbo-0125 setting1 0.12 0.12 33.10 0.46 6.02 7.96
DeepSeek-R1-32B setting1 13.27 13.32 13.69 13.55 13.50 13.47
DeepSeek-R1-32B setting2 16.71 16.74 16.62 16.71 16.71 16.70
DeepSeek-R1-32B setting4 21.20 21.39 21.35 21.39 21.39 21.34
DeepSeek-R1-32B setting5 39.37 40.57 39.94 40.57 40.57 40.20

### v3

@@
 This latter property we generally refer to as discriminant validity Cronbach (1951); Campbell & Fiske (1959).
-Our chosen measure is a simple aggregate of internal consistency and discriminant validity: Cronbach’s $\alpha$ for the former, and a cross-loading ratio (CLR) with sigmoid normalization centered at 1.5, and HTMT for the latter, formally –
-$R_{\text{psychometric}}$ $=\tfrac{1}{3}\Big(\tfrac{1}{k}\sum_{j=1}^{k}\alpha_{j}\;+\tfrac{1}{k}\sum_{j=1}^{k}\text{CLR}_{\text{norm},j}\;+\big(1-\tfrac{2}{k(k-1)}\sum_{i<j}\text{HTMT}_{ij}\big)\Big),$ (4)
-with sensitivity $\text{Sens}_{\text{psychometric}}=\sqrt{1-R_{\text{psychometric}}}\times\text{score\_range}$. Full definitions are provided in Section B.2.
+We aggregate internal consistency and discriminant validity using normalized, thresholded components and an explicit penalty for judge failure rates.
+Let $\alpha_{i}$ denote Cronbach’s alpha for factor $i$, $\text{CLR}_{i}$ its cross-loading ratio, $\overline{\text{HTMT}}_{i\cdot}$ the mean HTMT between factor $i$ and all others (computed on absolute item correlations), and $\phi_{i}\in[0,1]$ the fraction of missing/unscorable judgments for factor $i$.
+Define
+$\alpha_{i}^{\mathrm{norm}}=\tfrac{\alpha_{i}-0.70}{0.95-0.70}$
+$\text{CLR}_{i}^{\mathrm{norm}}=\tfrac{\text{CLR}_{i}-1.0}{2.0-1.0}$
+$\text{HTMT}i^{\mathrm{norm}}=1-\tfrac{1}{0.85}$
+Let $D_{i}$ be the harmonic mean of the two discriminant components, $D_{i}=\tfrac{2\,\text{CLR}^{\mathrm{norm}}_{i}\,\text{HTMT}^{\mathrm{norm}}_{i}}{\text{CLR}^{\mathrm{norm}}_{i}+\text{HTMT}^{\mathrm{norm}}_{i}}$.
+The per-factor validity is $V_{i}=\tfrac{1}{2}\alpha^{\mathrm{norm}}_{i}+\tfrac{1}{2}D_{i}$ and we apply a failure penalty multiplicatively, $\tilde{V}_{i}=(1-\phi_{i})V_{i}$.
+Our overall psychometric validity is the mean across factors,
+$R_{\text{psychometric}}=\tfrac{1}{k}\sum_{i=1}^{k}\tilde{V}_{i},$ (4)
+with sensitivity $\text{Sens}_{\text{psychometric}}=\sqrt{1-R_{\text{psychometric}}}\times\text{score\_range}$. Full definitions appear in Section B.2.
 3 Experimental Setup
@@
 Loadings heatmaps corroborate this collapse with extensive cross-loading and weak separation between intended constructs (Figure 5).
+Psychometric validity is effective for distinguishing quality judgment components from problematic ones.
+Using the validity formulation in Section B.2, we reevaluated Setting 1 across four judges (DeepSeek-R1-32B, GPT-3.5-Turbo-0125, GPT-4o-mini-0718, QwQ-32B).
+Figure 2 summarizes internal consistency (Cronbach’s $\alpha$), discriminant validity (HTMT computed on absolute correlations; CLR normalized to $[0,1]$ around a 1.0/2.0 baseline/target), and the resulting overall validity score (harmonic mean of discriminant subcomponents blended with $\alpha$, penalized by the factor-wise failure rate).
+Completeness, correctness, and style attain excellent internal consistency across judges ($\alpha\gtrsim 0.90$) and moderate discriminant signal.
+This suggests both that the Arena-Hard-Auto benchmark may be suitable for judging such factors, but also that judges frequently confuse them.
+Safety is systematically weaker: internal consistency is lower for some judges and the factor shows elevated failure rates (fraction of missing/unscorable judgments), which directly reduces validity.
+The most extreme outlier is GPT-3.5 on safety (largely due to failures of the judge to respond); this drives down the validity score despite otherwise reasonable separation on the remaining factors.
+Conciseness shows high validity – unfortunately, as we know from factor analysis, this is in part because it is ignored in the overall judgment.
+Figure 2: Psychometric validity summary for Setting 1 across four judges.
+Bars show (top-left) Cronbach’s $\alpha$ (internal consistency); (top-right) cross-loading ratio (CLR; higher indicates stronger factor separation, shown also as a normalized $[0,1]$ score in the validity computation); (bottom-left) HTMT computed on absolute item-level correlations (lower is better; the dashed line marks the 0.85 threshold); and (bottom-right) overall validity score combining $\alpha$ and discriminant components via a harmonic mean and applying a multiplicative penalty $(1-\phi)$ for factor-wise failure rate $\phi$ (share of unscorable judgments such as “Safety: N/A”).
+Atypical results appear primarily on the safety factor for GPT-3.5, where high failure rates depress the overall score.
 Figure 3: Most factors are highly correlated for most judges.
@@
 This suggests factor collapse – the inability of judges to meaningfully distinguish between semantically distinct rubric factors in the setting.
-Figure 3: Diverse judges exhibit very similar latent factor loadings.
-Across four different LLM judges in benchmark Setting 2, the eigenvalues associated with factor weightings are highly similar; all indicate a collapse of significance in the latent loadings.
 […]
@@
 Figure 4: ELO-style aggregation compresses multi-dimensional, noisy judgments into apparently smooth rankings, masking upstream uncertainty.
+Even more so than in general-case factor analysis, the feature loadings under ELO exhibit clear and strong positive or negative correlation.
 Ablations.
@@
 Table 1: Comprehensive analysis of explained variance across judge models and configurations. Open-source reasoning models show dramatically higher unexplained variance than closed-source models.
+Figure 5: Diverse judges exhibit very similar latent factor loadings.
+Across four different LLM judges in benchmark Setting 2, the eigenvalues associated with factor weightings are highly similar; all indicate a collapse of significance in the latent loadings.
 Appendix B Metrics
@@
   • Cross-loading Ratio (CLR): Quantifies factor discriminant validity
-  • Heterotrait-Monotrait Ratio (HTMT): Assesses construct validity between factors
+  • Heterotrait-Monotrait Ratio (HTMT): Assesses construct validity between factors (computed on absolute item–item correlations)
 Mathematical Formalization
@@
 Definition 8 (HTMT Ratio).
-The HTMT ratio between factors $f_{i}$ and $f_{j}$ is: $\text{HTMT}_{ij}=\frac{\bar{r}_{ij}}{\sqrt{\bar{r}_{ii}\cdot\bar{r}_{jj}}}$ where $\bar{r}_{ij}$ is the mean correlation between items of different factors, and $\bar{r}_{ii}$ is the mean correlation between items within factor $i$.
-Definition 9 (Sigmoid CLR Normalization).
-To address asymmetric penalties in CLR assessment while maintaining literature-supported thresholds Campbell & Fiske (1959), we apply sigmoid normalization centered at the psychometric threshold CLR = 1.5 from the 0.40-0.30-0.20 rule Hair et al.
-(2019), which states that items should have strong factor loadings on their intended factors and relatively low loadings on all secondary factors: $\text{CLR}_{\text{norm},i}=\frac{1}{1+\exp(-2(\text{CLR}_{i}-1.5))}$ This transformation provides smooth, bounded scoring where CLR = 1.5 yields 0.5, poor discriminant validity (CLR $\approx$ 0) yields $\approx$ 0.05, and strong validity (CLR $\geq$ 2.0) yields $\geq$ 0.73.
+The HTMT ratio between factors $f_{i}$ and $f_{j}$ is:
+$\text{HTMT}_{ij}=\frac{\overline{|r|}_{ij}}{\sqrt{\overline{|r|}_{ii}\cdot\overline{|r|}_{jj}}}$
+where $\overline{|r|}_{ij}$ is the mean absolute correlation between items of different factors, and $\overline{|r|}_{ii}$ is the mean absolute correlation between items within factor $i$.
+Using absolute correlations matches the conventional HTMT definition and avoids sign artefacts.
+Definition 9 (Bounded CLR Normalization).
+We map the cross-loading ratio to $[0,1]$ with a simple bounded linear transform using 1.0 as the “no separation” baseline and 2.0 as a strong separation target supported by practice Hair et al.
+(2019):
+$\text{CLR}_{\text{norm},i}=\mathrm{clip}\big((\text{CLR}_{i}-1.0)/(2.0-1.0),\,0,\,1\big).$
+This avoids asymmetric penalties and makes the scale interpretable: CLR $\leq$ 1.0 $\Rightarrow$ 0, CLR $\geq$ 2.0 $\Rightarrow$ 1.
+Definition 10 (Failure Rate).
+For each factor $f_{i}$, define a failure rate $\phi_{i}\in[0,1]$ as the fraction of judgments that are missing or unscorable for that factor (e.g., rubric omissions such as “Safety: N/A”).
+This term down-weights validity when a factor is rarely applied by a judge.
 Definition 11 (Unified Psychometric Validity Score).
-We combine these metrics into a unified psychometric validity score using equal weighting:
-$R_{\text{psychometric}}=\frac{1}{3}\left(\frac{1}{k}\sum_{i=1}^{k}\alpha_{i}+\frac{1}{k}\sum_{i=1}^{k}\text{CLR}_{\text{norm},i}+\left(1-\frac{2}{k(k-1)}\sum_{i<j}\text{HTMT}_{ij}\right)\right)$
-where the HTMT component is inverted since lower HTMT values indicate better discriminant validity.
+Let $\mathrm{clamp}(x;a,b)=\min\{\max\{x,a\},b\}$ and $H(a,b)=\tfrac{2ab}{a+b}$ denote the harmonic mean. Define per-factor components
+$\alpha^{\mathrm{norm}}_{i}$ $=\mathrm{clamp}\Big(\frac{\alpha_{i}-0.70}{0.95-0.70},\,0,\,1\Big),$
+$\text{CLR}^{\mathrm{norm}}_{i}$ $=\mathrm{clamp}\Big(\frac{\text{CLR}_{i}-1.0}{2.0-1.0},\,0,\,1\Big),$
+$\text{HTMT}^{\mathrm{norm}}_{i}$ $=\mathrm{clamp}\Big(1-\frac{1}{0.85}\cdot\overline{\text{HTMT}}_{i\cdot},\,0,\,1\Big),$
+where $\overline{\text{HTMT}}_{i\cdot}$ is the average HTMT between factor $i$ and all others. The discriminant component is
+$D_{i}=H\big(\text{CLR}^{\mathrm{norm}}_{i},\,\text{HTMT}^{\mathrm{norm}}_{i}\big).$
+The per-factor validity is then
+$V_{i}=\tfrac{1}{2}\,\alpha^{\mathrm{norm}}_{i}+\tfrac{1}{2}\,D_{i},$
+with a failure penalty applied multiplicatively: $\tilde{V}_{i}=(1-\phi_{i})\,V_{i}$. The overall psychometric validity aggregates across factors:
+$R_{\text{psychometric}}=\frac{1}{k}\sum_{i=1}^{k}\tilde{V}_{i}.$
 The psychometric reliability sensitivity, representing the unexplained variance due to poor factor structure and judge consistency, is then: $\text{Sensitivity}_{\text{psychometric}}=\sqrt{1-R_{\text{psychometric}}}\times\text{score\_range}$
@@
 Methodological Justification
-The sigmoid CLR normalization addresses several methodological concerns identified in our analysis:
+The bounded CLR normalization and the redundancy-aware discriminant component address several methodological concerns:
   • Literature grounding: The 1.5 threshold derives from established discriminant validity criteria where primary loadings $\geq$ 0.40 and cross-loadings ¡ 0.30 yield CLR $\approx$ 1.33-1.5 Hair et al.
 (2019)
-  • Balanced weighting: Prevents poor CLR values from creating excessive negative penalties that dominate well-established metrics (Cronbach’s $\alpha$, HTMT)
+  • Balanced weighting: Uses the harmonic mean of CLR and HTMT normalizations so that disagreement between the two signals is penalized rather than double-counted
   • Psychometric validity: Maintains standard practice in discriminant validity assessment while avoiding the asymmetric penalty problem of linear normalization Rönkkö & Cho (2022)
+B.3 Reliability Heatmaps, Failure Rates, and Distributions
+Figure 6: Internal consistency (Cronbach’s $\alpha$) by judge and factor (Setting 1).
+Higher is better; values near 1 indicate highly consistent item-level judgments within a factor.
+Most factors attain $\alpha\geq 0.9$ across judges; safety is comparatively weaker for some judges.
+Figure 7: Construct validity via HTMT (computed on absolute item correlations) by judge and factor.
+Lower is better; a conventional 0.85 threshold marks acceptable discriminant validity.
+Values near or above the threshold indicate factor overlap.
+Figure 8: Failure rates $\phi$ (fraction of missing/unscorable judgments) by judge and factor.
+This term down-weights validity: $\tilde{V}=(1-\phi)V$.
+The highest rates occur for GPT-3.5 on safety, reflecting frequent rubric omissions (e.g., “Safety: N/A”).
+Figure 9: Distribution of reliability/validity components across factors.
+Violin plots summarize dispersion of Cronbach’s $\alpha$, CLR, HTMT, and the composite validity score.
+Narrow, high-location violins indicate strong, stable performance; wide, tall, or low-location violins flag potential instability.
 Appendix C Experimental Evaluation Settings